Sami0627:

Jason has two bags with 6 tiles each. The tiles in each bag are shown below: Make 6 squares. The squares are numbered sequentially from 1 to 6. Without looking, Jason draws a tile from the first bag and then a tile from the second bag. What is the probability of Jason drawing an even tile from the first bag and an even tile from the second bag? 6 over 12 9 over 12 6 over 36 9 over 36

Angle:

Because this is an "AND" question, we will be multiplying two probabilities We want: (prob of even tile from first bag) x (prob of even tile from second bag) = (combined prob)

Angle:

what's the probability of getting an even tile from the first bag?

Sami0627:

the probability is 1/6?

Angle:

hmmm out of 1, 2, 3, 4, 5, 6 what is the chance that I get an even number?

Sami0627:

3%?

Angle:

how'd you get 3%?

Angle:

2, 4, and 6 are even so the chance of getting an even number out of 1, 2, 3, 4, 5, 6 would be. 3 out of 6

Sami0627:

because I count like 1 is odd 2 is even. so what I think 2 4 and 6 are even

Angle:

ok but then we will have the fraction \(\large \frac{3}{6}\) because there are 6 numbers there total, so it won't be 3 percent, it would be 3 out of 6

Sami0627:

oh ok

Angle:

it would be the same probability for the second bag right?

Sami0627:

yes

Angle:

Then (prob of even tile from first bag) x (prob of even tile from second bag) = (combined prob) which means: \(\large \frac{3}{6}\times\frac{3}{6} =\) ?

Sami0627:

9/36?

Angle:

yup!

Sami0627:

thank you for your help